OpenStudy (anonymous):

will give a medal!!! What is the solution to the system of equations? x = 4y + 6 x - 3y = 4

OpenStudy (alexandervonhumboldt2):

4+3y=4y+6 -y=2 y=-2 x=-2

OpenStudy (alexandervonhumboldt2):

(-2, -2)

OpenStudy (anonymous):

substitution (4y + 6) - 3y = 4 4y + 6 -3y = 4 (4y -3y) + 6 = 4 y + 6 - 6 = 4 - 6 y = -2 x = 4y + 6 x = -8 + 6 x = -2 (-2, -2)

OpenStudy (alexandervonhumboldt2):

there are two methond but both are correect